# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

### v2

@@ A.8 Detailed Result Table
Table 9: Machine Translation Results for English to German on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93
Base 37.17 18.57 87.13
15 Skip Decode Router + LoRA 23.04 10.52 55.62
Router 3.99 1.18 23.33
Random Skip Router + LoRA 30.43 10.98 66.25
Router 26.54 8.77 60.27
Unified Skip Router + LoRA 28.92 10.64 59.34
Router 23.23 7.85 59.26
FiRST (Ours) Router + LoRA 38.01 17.89 82.14
Router 28.83 11.8 67.74
25 Skip Decode Router + LoRA 13.67 6.00 31.47
Router 3.24 0.92 21.55
Random Skip Router + LoRA 6.01 0.91 29.71
Router 3.65 0.49 29.95
Unified Skip Router + LoRA 15.67 3.36 31.69
Router 12.58 2.65 32.15
FiRST (Ours) Router + LoRA 17.84 4.14 34.95
Router 9.67 1.37 26.01
35 Skip Decode Router + LoRA 5.55 0.33 23.85
Router 3.03 0.82 20.03
Random Skip Router + LoRA 1.80 0.12 25.56
Router 1.44 0.06 25.34
Unified Skip Router + LoRA 6.44 0.77 22.05
Router 3.92 0.51 22.88
FiRST (Ours) Router + LoRA 6.39 0.42 19.96
Router 3.7 0.14 21.41
Table 10: Machine Translation Results for English to German on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 51.81 30.04 79.13
Base 32.10 17.92 61.84
15 Skip Decode Router + LoRA 38.14 21.68 46.70
Router 9.59 4.84 34.14
Random Skip Router + LoRA 38.75 17.39 57.79
Router 13.56 5.64 35.86
Unified Skip Router + LoRA 36.96 17.35 57.10
Router 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 45.66 23.66 67.45
Router 22.69 10.92 54.55
25 Skip Decode Router + LoRA 27.81 15.74 42.01
Router 7.01 3.41 29.51
Random Skip Router + LoRA 25.12 9.85 45.53
Router 10.07 3.69 31.50
Unified Skip Router + LoRA 30.56 12.27 42.30
Router 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 32.92 13.74 41.66
Router 10.46 3.55 27.83
35 Skip Decode Router + LoRA 2.84 1.10 22.30
Router 2.36 1.04 18.32
Random Skip Router + LoRA 4.17 1.52 32.81
Router 2.11 0.76 23.63
Unified Skip Router + LoRA 3.26 0.52 19.12
Router 2.20 0.37 19.82
FiRST (Ours) Router + LoRA 20.01 6.96 28.10
Router 6.54 1.78 23.13
@@
Table 11: Machine Translation Results for English to Chinese on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 56.94 35.56 82.66
Base 38.02 22.46 68.95
15 Skip Decode Router + LoRA 28.73 15.84 55.98
Router 4.74 2.33 21.75
Random Skip Router + LoRA 47.88 25.75 67.32
Router 36.60 18.66 59.89
Unified Skip Router + LoRA 46.61 25.01 69.58
Router 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 48.35 26.57 68.63
Router 17.55 8.68 42.76
25 Skip Decode Router + LoRA 20.03 10.85 33.85
Router 3.78 1.84 20.93
Random Skip Router + LoRA 11.69 4.66 27.73
Router 7.37 2.81 35.16
Unified Skip Router + LoRA 34.90 15.75 50.59
Router 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 35.79 15.66 56.92
Router 11.01 3.23 25.45
35 Skip Decode Router + LoRA 12.24 6.27 25.23
Router 3.64 1.74 22.84
Random Skip Router + LoRA 7.38 2.19 27.35
Router 4.47 1.12 29.46
Unified Skip Router + LoRA 7.51 2.10 20.25
Router 3.87 1.06 21.24
FiRST (Ours) Router + LoRA 15.66 3.95 26.80
Router 6.13 1.54 22.89
Table 12: Machine Translation Results for English to Chinese on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
@@
Layer $\downarrow$ R R+L R R+L R R+L
$\alpha\rightarrow$ 0.005 0.005 0.01 0.01 0.025 0.025
0 0.00 0.00 0.00 0.00 0.00 0.00
1 0.00 0.00 0.00 0.00 0.00 0.00
2 0.00 0.00 0.00 0.00 0.00 0.00
3 0.00 0.00 0.00 0.00 0.00 0.00
4 0.00 0.00 0.00 0.00 0.00 0.00
5 0.00 0.00 0.00 0.00 0.00 0.00
6 0.00 0.00 0.00 0.00 0.00 0.00
7 100.00 100.00 100.00 100.00 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00
9 100.00 100.00 0.00 0.00 0.10 1.32
10 0.00 0.00 0.00 0.00 89.25 72.67
11 0.00 0.00 0.00 0.00 0.00 0.00
12 0.00 0.00 100.00 100.00 100.00 100.00
13 0.00 0.00 0.00 0.00 0.00 0.00
14 0.00 0.00 0.00 0.59 0.00 0.25
15 0.00 0.00 100.00 99.95 100.00 99.36
16 0.00 0.00 0.10 2.45 100.00 99.95
17 0.00 0.00 0.00 0.00 0.00 0.00
18 97.79 32.24 100.00 100.00 100.00 100.00
19 0.00 0.00 99.85 91.17 100.00 99.61
20 0.00 0.00 100.00 100.00 100.00 99.46
21 99.85 98.72 100.00 100.00 100.00 100.00
22 0.00 0.00 0.00 0.00 0.00 0.00
23 0.00 0.00 24.14 0.79 99.75 91.66
24 0.00 0.00 0.00 0.00 0.00 0.00
25 0.00 0.00 0.00 0.00 0.00 0.00
26 0.00 0.00 57.31 2.45 100.00 86.02
27 0.00 0.00 0.00 0.00 0.00 0.00
28 0.00 0.00 0.00 0.00 0.00 0.05
29 0.00 0.00 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00 0.00 0.00
31 0.00 0.00 0.00 0.00 0.00 0.00
Avg 15.55 13.47 27.54 24.92 37.16 35.95
Table 15: Variation in skipping percentage (15-35%) with the Non-skip Penalization Loss coefficient $\alpha$ for LLaMA-3-8B on Machine Translation (English-to-German). As $\alpha$ increases, the skipping percentage also increases for both the Router-Only and Router-Augmented LoRA fine-tuned models. Similar trends are observed for other datasets as well.

### v3

@@ A.6 Layer-wise Skipping Statistics
Table 6: Machine Translation Results for English to German on LLaMA-3.2-3B and LLaMA-3-8B: BLEU and COMET scores for various skipping strategies.
Skip (%) Model Type Config BLEU-1 BLEU-2 BLEU COMET
LLaMA-3.2-3B
0 Original Model Base + LoRA 0.518 0.300 0.278 79.13
Base 0.321 0.179 0.166 61.84
15 Skip Decode Router + LoRA 0.381 0.217 0.200 46.70
Router 0.096 0.048 0.062 34.14
Random Skip Router + LoRA 0.387 0.174 0.154 57.79
Router 0.136 0.056 0.077 35.86
Unified Skip Router + LoRA 0.370 0.173 0.157 57.10
Router 0.224 0.094 0.087 45.16
FiRST (Ours) Router + LoRA 0.457 0.237 0.218 67.45
Router 0.227 0.109 0.100 54.55
25 Skip Decode Router + LoRA 0.278 0.157 0.147 42.01
Router 0.070 0.034 0.045 29.51
Random Skip Router + LoRA 0.251 0.099 0.137 45.53
Router 0.101 0.037 0.053 31.50
Unified Skip Router + LoRA 0.306 0.123 0.108 42.30
Router 0.152 0.055 0.079 31.63
FiRST (Ours) Router + LoRA 0.329 0.137 0.126 41.66
Router 0.105 0.035 0.053 27.83
35 Skip Decode Router + LoRA 0.028 0.011 0.016 22.30
Router 0.024 0.010 0.014 18.32
Random Skip Router + LoRA 0.042 0.015 0.022 32.81
Router 0.021 0.008 0.011 23.63
Unified Skip Router + LoRA 0.033 0.005 0.010 19.12
Router 0.022 0.004 0.007 19.82
FiRST (Ours) Router + LoRA 0.100 0.070 0.075 28.10
Router 0.065 0.018 0.029 23.13
LLaMA-3-8B
0 Original Model Base + LoRA 0.569 0.356 0.333 82.66
Base 0.380 0.225 0.208 68.95
15 Skip Decode Router + LoRA 0.287 0.158 0.149 55.98
Router 0.047 0.023 0.030 21.75
Random Skip Router + LoRA 0.479 0.258 0.237 67.32
Router 0.366 0.187 0.168 59.89
Unified Skip Router + LoRA 0.466 0.250 0.230 69.58
Router 0.273 0.134 0.122 54.57
FiRST (Ours) Router + LoRA 0.484 0.266 0.247 68.63
Router 0.176 0.087 0.080 42.76
25 Skip Decode Router + LoRA 0.173 0.094 0.088 33.85
Router 0.038 0.018 0.024 20.93
Random Skip Router + LoRA 0.117 0.047 0.065 27.73
Router 0.074 0.028 0.040 35.16
Unified Skip Router + LoRA 0.349 0.158 0.142 50.59
Router 0.177 0.074 0.068 38.74
FiRST (Ours) Router + LoRA 0.358 0.157 0.119 56.92
Router 0.110 0.032 0.051 25.45
35 Skip Decode Router + LoRA 0.112 0.057 0.054 25.23
Router 0.036 0.017 0.023 22.84
Random Skip Router + LoRA 0.074 0.022 0.034 27.35
Router 0.045 0.011 0.018 29.46
Unified Skip Router + LoRA 0.075 0.021 0.034 20.25
Router 0.039 0.011 0.017 21.24
FiRST (Ours) Router + LoRA 0.157 0.040 0.066 26.80
Router 0.061 0.015 0.025 22.89
@@
Table 7: Machine Translation Results for English to Chinese on LLaMA-3.2-3B and LLaMA-3-8B: BLEU and COMET scores for various skipping strategies.
Layer $\downarrow$ $\alpha=0.005$ $\alpha=0.01$ $\alpha=0.025$
R R+L R R+L R R+L
0 0.00 0.00 0.00 0.00 0.00 0.00
1 0.00 0.00 0.00 0.00 0.00 0.00
2 0.00 0.00 0.00 0.00 0.00 0.00
3 0.00 0.00 0.00 0.00 0.00 0.00
4 0.00 0.00 0.00 0.00 0.00 0.00
5 0.00 0.00 0.00 0.00 0.00 0.00
6 0.00 0.00 0.00 0.00 0.00 0.00
7 100.00 100.00 100.00 100.00 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00
9 100.00 100.00 0.00 0.00 0.10 1.32
10 0.00 0.00 0.00 0.00 89.25 72.67
11 0.00 0.00 0.00 0.00 0.00 0.00
12 0.00 0.00 100.00 100.00 100.00 100.00
13 0.00 0.00 0.00 0.00 0.00 0.00
14 0.00 0.00 0.00 0.59 0.00 0.25
15 0.00 0.00 100.00 99.95 100.00 99.36
16 0.00 0.00 0.10 2.45 100.00 99.95
17 0.00 0.00 0.00 0.00 0.00 0.00
18 97.79 32.24 100.00 100.00 100.00 100.00
19 0.00 0.00 99.85 91.17 100.00 99.61
20 0.00 0.00 100.00 100.00 100.00 99.46
21 99.85 98.72 100.00 100.00 100.00 100.00
22 0.00 0.00 0.00 0.00 0.00 0.00
23 0.00 0.00 24.14 0.79 99.75 91.66
24 0.00 0.00 0.00 0.00 0.00 0.00
25 0.00 0.00 0.00 0.00 0.00 0.00
26 0.00 0.00 57.31 2.45 100.00 86.02
27 0.00 0.00 0.00 0.00 0.00 0.00
28 0.00 0.00 0.00 0.00 0.00 0.05
29 0.00 0.00 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00 0.00 0.00
31 0.00 0.00 0.00 0.00 0.00 0.00
Avg 15.55 13.47 27.54 24.92 37.16 35.95
Table 8: Variation in skipping percentage (15–35%) with the non-skip penalization loss coefficient $\alpha$ for LLaMA-3-8B on English–German translation. As $\alpha$ increases, average skipping rises across both Router-only (R) and Router+LoRA (R+L) models.
@@
[…]
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.88x
R + LoRA 25 0.82x 0.83x
R + LoRA 35 0.69x 0.68x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 20 0.81x
R + LoRA 27 0.76x
Table 11: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST. The reported values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
[…]
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.91x
R + LoRA 25 0.78x 0.75x
R + LoRA 35 0.69x 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 24 0.79x
R + LoRA 28 0.77x
Table 12: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST. The reported values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
A.7 Detailed Result Table
Tables 6 and 7 present the detailed results for the LLaMA-3.2-3B and LLaMA-3-8B models for an additional skipping percentage on Machine Translation Task. The results are reported using BLEU (BLEU-1, BLEU-2, BLEU) and COMET metrics, highlighting performance across different skipping percentages. Similarly, Table 9 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L. Lastly, Table 10 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
A.8 Generalizability and Reusability of Routers
We also investigate whether a router trained on one dataset can be applied directly to another dataset for the same task, without any additional training. Analysis of layer-skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere. To validate this idea, we train routers on English→German translation data and test them on English→Chinese translation, and likewise train on SQuAD for question answering before evaluating on the Natural Questions dataset. The quantitative results of these cross-dataset evaluations are presented in Table 13. In the QA experiments, the router trained and tested on NQ itself (denoted as Original in Table 13) and the router trained on SQuAD but tested on NQ (Cross-dataset) both skip a comparable fraction of layers. Despite being trained on different data, the SQuAD-trained router achieves QA accuracy (EM score) very close to that of the NQ-trained router, as detailed in Table 13. This indicates that the essential decision patterns learned by the router transfer well across QA datasets when using a similar level of layer skipping. A similar pattern emerges in machine translation. Whether trained on English→German or directly on English→Chinese data, routers that skip the same proportion of layers produce very similar translation quality metrics (BLEU and COMET in Table 13) on the English→Chinese task. This holds across different model sizes and skip settings. Together, these results demonstrate that routers learned on one dataset can be effectively reused on another dataset for the same task, as long as they are configured to skip a similar amount of computation. This reusability can lead to substantial savings in both training time and computing resources.
Question Answering (Natural Questions)
Setting Skip (%) EM
LLaMA-3.2-3B
Original 10.32 22.55
Cross-Dataset 12.46 20.87
Original 18.03 17.51
Cross-Dataset 19.04 16.93
LLaMA-3-8B
Original 11.44 33.53
Cross-Dataset 13.11 29.70
Original 24.60 18.02
Cross-Dataset 26.80 19.58
Machine Translation (Eng–Zh)
Setting Skip (%) BLEU COMET
LLaMA-3.2-3B
Original 14.69 0.37 0.55
Cross-Dataset 16.57 0.37 0.45
Original 26.12 0.25 0.28
Cross-Dataset 26.16 0.25 0.30
LLaMA-3-8B
Original 16.46 0.25 0.43
Cross-Dataset 14.70 0.37 0.49
Original 27.73 0.25 0.25
Cross-Dataset 30.00 0.25 0.24
[…]